OpenStudy (anonymous):

During the first part of a trip a canoeist travels 54 miles at a certain speed. The canoeist travels 20 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 4 hours. What was the speed on each part of the trip?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

d=r*t (d is distance, r is speed, t is time) First part: 54 = r*t1 second part: 20 = (r-5)*t2 t1 + t2 = 4 system of 3 equations with 3 variables use substitution t1 = 4-t2 r = 54/t1 => t1 = 54/r 54/r = 4-t2 => t2 = 4-54/r \[20 = (r-5)(4-\frac{54}{r})\] \[20 = 4r +\frac{270}{r} -54-20\] \[4r -94 + \frac{270}{r} = 0\] multiply everything by r \[4r^{2}-94r +270 = 0\] quadratic formula \[r = \frac{94 \pm 67.2}{8}\] \[r = 3.35 , 20.15\] since you cant go neg speed, r>5 thus r=20.15 2 speeds: 20.15 and 15.15

OpenStudy (dumbcow):

were you able to follow and understand how i got the answers?

OpenStudy (anonymous):

yes I was thank you

OpenStudy (dumbcow):

:)